a)        Air enters the compressor of a gas turbine at 100 kPa and 25°C. For a pressure ratio of 5 and a maximum temperature of 850°C determine the back work ratio and the thermal efficiency using the Brayton cycle.



1
Expert's answer
2021-09-21T02:06:42-0400



Given information;

Initial pressure, "p_1=100kPa"

Initial temperature, "T_1=25\u00b0 C"

Compression ratio, r =5

Turbine inlet temperature, "T_3=850\u00b0C"

At temperature "T_1=25\u00b0C=298 K"

and the pressure "p_1=100 kPa"


"h_1=295.17+{(298-295)\/(300-295)}*(300.9-295.17)=298.608 kJ\/kg"


"(p_r)_1=1.3068+{(298-295)\/(300-295)}*(1.386-1.3068)=1.35432"


"(p_r)_2=(p_2\/p_1)((p_r)_1= 5 \u00d71.35432=6.7716"


At "(p_r)_2=6.7716"

By interpolating

"h_2=472.24+{(6.7716-6.742)\/(7.268-6.742)}*(482.49-472.24)=472.8168 kJ\/kg"


At temperature "T_3=850\u00b0 C=1123 K"

"h_3=1184.28+{(1123-1120)\/(1140-1120)}\u00d7(1207.57-1184.28)=1187.77 kJ\/kg"


"(p_r)_3=179.7+{(1123-1120)\/(1140-1120)}\u00d7(193.1-179.7)=181.71"


"(p_r)_4=(p_4\/p_3)((p_r)_3= 1\/5 \u00d7181.71=36.342"



At "(p_r)_4" = 36.342

By interpolation

"h_4=756.44+{(36.342-35.5)\/(37.35-35.5)}\u00d7(767.29-756.44)=767.38 kJ\/kg"


We know the thermal efficiency of the cycle

"η=\dfrac{(1187.77-767.38)-(472.8168-298.608)}{1187.77-472.8168}"


"η=0.344\\



η=34.4 \%"


The back work ratio

"bwr = W_c\/W_t=(h_2-h_1)\/(h_3-h_4)"


"bwr =\dfrac{(472.8168-298.608)}{(1187.77-767.38)}"


"bwr =0.4144\\


bwr = 41.44 \%"